OpenStudy (anonymous):

derivative of (ln x)/4

Directrix (directrix):

Wouldn't this be [1/(4x)] ? (1/4) derivative of ln(x) = (1/4) (1/x) = 1/(4x)

OpenStudy (anonymous):

It is 1/4x and for all of questions like this just go to http://www.wolframalpha.com/ if you don't know things

Directrix (directrix):

It's not where I went to learn mathematics but I do find it to be a helpful tool. And, one of many tools I use.

OpenStudy (anonymous):

derivative of (ln x)/4 derivative of lnx is 1/x derivative of (ln x)/4 is 1/4x.